OpenStudy (anonymous):

Find the correspoding x and y equation for the curve: x=t^2 -1 y=t^2 +1 my solution : x + 1= t^2 so t = sqrt of x+1 y-1 = t ^2 , so t = sqrt of y-1 is it correct ? thanks in advance.

OpenStudy (shiraz14):

x = y-2 y = x+2